± pH Changes in Buffers

Part A

What is the pH of a buffer prepared by adding 0.708 mol of the weak acid HA to 0.608 mol of NaA in 2.00 L of solution? The dissociation constant Ka of HA is 5.66×10−7.

Express the pH numerically to three decimal places.

Part B

What is the pH after 0.150 mol of HCl is added to the buffer from Part A? Assume no volume change on the addition of the acid.

Express the pH numerically to three decimal places.

Part C

What is the pH after 0.195 mol of NaOH is added to the buffer from Part A? Assume no volume change on the addition of the base.

Express the pH numerically to three decimal places.

View Available Hint(s)
Answer #1

part A

pH of acidic buffer = pka + log(NaA/HA)

    pka =- log(5.66*10^-7) = 6.25

pH = 6.25+log(0.608/0.708)

   = 6.184

part B

No of mol Of HCl = 0.15 mol

pH of acidic buffer = pka + log((NaA-HCl)/(HA+HCl))

            = 6.25+log((0.608-0.15)/(0.708+0.15))

            = 5.977

part C


No of mol Of NaOH = 0.195 mol

pH of acidic buffer = pka + log((NaA+NaOH)/(HA-NaOH))

            = 6.25+log((0.608+0.195)/(0.708-0.195))

             = 6.444